OpenStudy (anonymous):

I am trying to input this question into my ti-84 plus calculator and it will not work. Is there a different way to solve this question? log4 (64k^4x) This is supposed to have an answer without exponents if that make a difference on how I do it.

OpenStudy (b87lar):

what is the question? Are you supposed to simply this expression?

OpenStudy (anonymous):

The expression is: log4 (64 k^4 x) Yes it is supposed to be simplified

OpenStudy (b87lar):

\[\log_4(64 k^4 x)=\log_4 64+\log_4 k^4 +\log_4 x=\] \[3\log_4 4+4\log_4 k+\log_4 x=\] \[3+4\log_4 k +\log_4 x\] that's probably the simplest form

OpenStudy (anonymous):

Okay thank you very much! I'm not doing very well with all of the log functions. Have a good night!